OpenStudy (anonymous):

Find the zeros of the function f(t) = 2t^3 + 3t^2 + 8t +12

OpenStudy (anonymous):

0=2t^3 + 3t^2 + 8t +12 0=(2t+3)(t^2+4) 0=2t+3 or 0=t^2+4 2t=-3 or t^2=-4 t=-3/2 and the square root of a negative number does not exist, so t=-3/2

OpenStudy (anonymous):

t^2=-4 has the solution t=2i or t=-2i